2/3x+3/5y=-17
1/2x-1/3y=-1
solve using elimination

Found 2 solutions by John10, stanbon:
Answer by John10(297)

2/3x+3/5y=-17
1/2x-1/3y=-1
Rewrite each equation by multiplying the LCD of each equation respectively
10x + 9y = -255
3x - 2y = -6
Use elimination method to solve this system.
Hint: the answer is x = -12 and y = -15
There you go!

Answer by stanbon(75887)

2/3x+3/5y=-17
1/2x-1/3y=-1
==============
Multiply thru 1st by 15.
Multiply thru 2nd by 6.
--------------
10x + 9y = -15*17
3x - 2y = -6
---
Multiply thru 1st by 3
Multiply thru 2nd by 10
-----
30x + 27y = -45*17
30x - 20y = -60
-----------------------
Subtract and solve for "y":
47y = -705
y = -15
---
Solve for "x":
3x-2y = -6
3x+30 = -6
3x = -36
x = -12
===========
Cheers,
Stan H.
===========